OpenStudy (anonymous):

The center of a circle is at (4, -1) and it has a radius of 6. What is the equation of the circle?

OpenStudy (hoblos):

the equation of a circle with center (a,b) and radius r is \[(x-a)^2 + (y-b)^2 = r^2\] can you apply this ?

OpenStudy (anonymous):

(4 - a) + (-1 + b)^2 =6^2 ?

OpenStudy (hoblos):



OpenStudy (hoblos):

keep the x and y

OpenStudy (anonymous):

(x + 4) + (y - 1 )^2 =36

OpenStudy (hoblos):

if a=4 , -a = ?

OpenStudy (anonymous):

so just -4 instead of positive?

OpenStudy (hoblos):

you have to check the -1 also

OpenStudy (anonymous):

and the -1 is a +1 ?

OpenStudy (hoblos):

correct